# Restoring Calibration for Aligned Large Language Models: A Calibration-Aware Fine-Tuning Approach

### v1

@@ 6.3 Calibration Plots
  • (d) RCFT (0.0897): While RCFT is not as well-calibrated as CFT, it still surpasses DPO and TS in bringing predicted probabilities in line with actual accuracies. Its con-ECE is 0.0897, representing a moderate level of calibration improvement. Note that RCFT has a better improvement on accuracy as indicated in Table 2.
@@
To conclude this section, we present an ablation study examining the effects of using $\mathcal{L}_{\text{SFT}_{2}}$ or $\mathcal{L}_{\text{ECE}}$ in isolation. Using only $\mathcal{L}_{\text{SFT}_{2}}$ achieves high accuracy ( 90%) but shows poor calibration. Conversely, using only $\mathcal{L}_{\text{ECE}}$ drives the model toward random guessing, achieving near-zero ECE but with only 25% accuracy. These results, detailed in Section C.2, demonstrate the importance of balancing both objectives to achieve both good calibration and performance.
@@ A.1 Preference Optimization
The original DPO method (Rafailov et al., 2023) is to directly optimize of the policy without explictly training the reward function in a supervised manner:
@@ A.2 Other Related Work
Several notable variants of DPO have been developed (Liu et al., 2023; Azar et al., 2024; Chang et al., 2024; Gorbatovski et al., 2024; Rafailov et al., 2024a; Yang et al., 2024). However, recent studies (Li et al., 2023a; Xu et al., 2024; Tajwar et al., 2024) suggest that DPO is less effective than reward-based RLHF methods for aligning LLMs. Both Li et al. (2023a) and Xu et al. (2024) attributed this shortcoming to representation misspecification in DPO, which limits its ability to achieve robust alignment compared to reinforcement learning approaches such as PPO. Additionally, the on-policy nature of reward-based fine-tuning helps mitigate distribution shifts between the training data and online responses, thereby enhancing LLM performance (Tajwar et al., 2024).
Nash learning from human feedback (NLHF) has also been proposed as a framework to align LLMs with general preference models (Munos et al., 2023; Liu et al., 2025). Moreover, Wang et al. (2025) analyzed the convergence behavior of NLHF, showing that it approaches the Nash equilibrium in the last iterate.
@@ A.4 Rank Preserving Mapping
We Denote $\text{Conf}_{\pi}(x)=[c_{1},c_{2},c_{3},c_{4}]$. Without loss of generality, let us assume that $c_{1}\geq c_{2}\geq c_{3}\geq c_{4}$. Let $\boldsymbol{p}(x)=[c^{\prime}_{1},c^{\prime}_{2},c^{\prime}_{3},c^{\prime}_{4}]$ be the target confidence. Firstly, we have $c^{\prime}_{1}=q_{m}$, the accuracy within the m-th bin.
For $i=2,3,4$, since we would like to adjust the max confidence to bin accuracy $q_{m}$, and meanwhile reserve the order of four confidences, we would have to force the new confidence distribution of the rest of three options be $1-q_{m}$. Hence we can apply a linear + non-linear transformation on them:
$c_{i}^{\prime}=\alpha\tanh{(\gamma c_{i})}+\beta.$
First, let us discuss how to find the parameter $\gamma$ for non-linear mapping. Note that a higher $\gamma$ will make $\tanh(\cdot)$ saturate to be closed to $1$. This is to measure the uncertainty level between the rest three option. We simply let
$\tanh(\gamma\max\{c_{2},c_{3},c_{4}\})=0.99.$
To avoid the non-invertible situation, we set it to be 0.99 rather than 1. Then, we have
$\gamma\max\{c_{2},c_{3},c_{4}\}=\text{artanh}(0.99)\approx\ln{(3)}$
$\gamma=\frac{\ln{(3)}}{\max\{c_{2},c_{3},c_{4}\}}\cdot\frac{1}{1-q_{m}},$
$\sum_{i=2}^{4}(\alpha\cdot\tanh{(\gamma c_{i})}+\beta)=1-q_{m},$
$\alpha\sum_{i=2}^{4}\tanh{(\gamma c_{i})}+3\beta=1-q_{m}.$
How do we get the desired $\alpha$ and $\beta$? Since we will adjust $c_{1}$ towards $q_{m}$, we need to make sure that $\sum_{i=2}^{4}c_{i}^{\prime}=1-q_{m}$, let $T=\sum_{i=2}^{4}\tanh{(\gamma c_{i})}.$ And hence
$\sum_{i=2}^{4}\alpha(\tanh{\gamma c_{i}}+\beta)$ $=1-q_{m},$
$\alpha T+3\beta=1-q_{m},$
$\alpha=\frac{1-q_{m}-3\beta}{T}.$
Assume that $\beta=k\alpha$. For simplicity, let $k=1$, hence $\beta=\alpha$. Then
$\alpha T+3\alpha=1-q_{m},$
$\alpha(T+3)=1-q_{m},$
$\alpha=\frac{1-q_{m}}{T+3}.$
For simplicity, we can estimate $\alpha$, and then solve for $\beta$ according to the value of $\alpha$:
$\alpha=\frac{1-q_{m}}{\sum_{i=2}^{4}\tanh{(\gamma c_{i})+3}}.$
And then find $\beta$ by:
$\beta=\frac{1-q_{m}-\alpha\sum_{i=2}^{4}\tanh{(\gamma c_{i})}}{3}.$
To check if $c_{i}^{\prime}<q_{m}$, we have
$\alpha\tanh{(\gamma c_{i})}+\beta$ $<q_{m}.$
Since $\tanh{(\cdot)}\in(-1,1)$, we have:
@@
Then for $\alpha,\beta$,
$\alpha<\frac{1-q_{m}}{3},\beta<\frac{1-q_{m}}{3}.$
Then, we can get
$\frac{2-q_{m}}{3}$ $<q_{m},$
$2-2q_{m}$ $<3q_{m},$
$q_{m}$ $>0.25.$
Hence, when $q_{m}>0.25$, the rank ordering of all four options can be preserved. When $q_{m}\leq 0.25$, the parameter $\gamma$ ensures that the non-linear mapping sufficiently compresses $c_{i}^{\prime}$ close to $1$. After applying the linear mapping with parameters $\alpha$ and $\beta$, we maintain both the rank ordering of the remaining three confidence scores and keep their differences small, as low accuracy implies high uncertainty, requiring similar confidence values across options.
@@ Appendix B Proof of Technical Results
Then, We define $\pi(x)$ as follow. When $x\in\mathcal{S}_{\textnormal{T}}$ or $\mathcal{S}_{\textnormal{F,2}}$,
@@
Then, We define $\pi(x)$ as follow. When $x\in\mathcal{S}_{\textnormal{F}}$ or $\mathcal{S}_{\textnormal{T,2}}$,
@@ C.1 Comprehensive Descriptions for Models, Baselines, Datasets, and Metrics
Models. In our study, we employ four widely-used open-source large language models to investigate the calibration issue and validate the effectiveness of our proposed method. They include
@@
  • Vicuna-7B-v1.5222https://huggingface.co/lmsys/vicuna-7b-v1.5: Vicuna-7B-v1.5 (Chiang et al., 2023) is a chat assistant developed by LMSYS, fine-tuned from Llama 2 using approximately 125,000 user-shared conversations collected from ShareGPT.com. This auto-regressive language model employs the transformer architecture and is designed for research purposes in natural language processing, machine learning, and artificial intelligence. The model has been evaluated using standard benchmarks, human preferences, and LLM-as-a-judge methodologies.
@@
Baselines We validate the effectiveness of the proposed method by comparing performance among four kinds of models.
@@
Dataset To evaluate the efficacy of our proposed calibration method, we employ five datasets to conduct comprehensive experiments:
@@
Metric The discrete conf-ECE and cw-ECE are defined as follows.
Confidence Expected Calibration Error (conf-ECE)
@@
where $B_{m}$ is the $j$-th bin in terms of max confidence; $|B_{m}|$ denotes the cardinality of the bin; $\mathbb{P}(Y=I|\mathbf{x}\in B_{m})$ and $\mathbb{E}[\pi_{\theta}(\mathbf{y}=I|\mathbf{x})|\mathbf{x}\in B_{m}]$ denote the average prediction of winning class probability and the actual probability that the winning class is the ground truth class. Class-Wise Expected Calibration Error (cw-ECE)
@@
We then introduce the detail of the Win Rate metric. Win Rate metric evaluates the model’s ability to distinguish between preferred and non-preferred responses within a given dataset. To this end, we utilize a widely-used preference dataset  [^10] https://huggingface.co/datasets/argilla/ultrafeedback-binarized-preferences, which provides pairs of responses: a chosen_response (preferred) and a reject_response (non-preferred) for each instruction. We then calculate the sequence probabilities of these provided responses under the target model’s distribution.
The sequence probability of a response $R=(w_{1},w_{2},\ldots,w_{T})$, where $w_{i}$ denotes the $i$-th token, is computed as:
@@
Here, $P(w_{i}\mid w_{1},w_{2},\ldots,w_{i-1})$ represents the probability of token $w_{i}$ given its preceding context, as determined by the model. For each instruction in the dataset, we calculate the sequence probabilities $P_{\text{chosen}}$ and $P_{\text{reject}}$ for the chosen_response and reject_response, respectively. A “win” is recorded if $P_{\text{chosen}}>P_{\text{reject}}$; otherwise, it is a “loss.” The Win Rate is then defined as the proportion of wins over the total number of comparisons:
Win Rate=Number of WinsTotal Number of $\text{Win Rate}=\frac{\text{Number of Wins}}{\text{Total Number of Comparisons}}$

### v2

@@ 6.3 Calibration Plots
  • (d) RCFT (0.0897): While RCFT is not as well-calibrated as CFT, it still surpasses DPO and TS in bringing predicted probabilities in line with actual accuracies. Its con-ECE is 0.0897, representing a moderate level of calibration improvement. Note that RCFT has a better accuracy as indicated in Table 2.
@@
To conclude this section, we present an ablation study examining the effects of $\lambda$. Using only $\mathcal{L}_{\text{SFT}_{2}}$ achieves high accuracy ( 90%) but shows poor calibration. Conversely, using only $\mathcal{L}_{\text{ECE}}$ drives the model toward random guessing, achieving near-zero ECE but with only 25% accuracy. These results, detailed in Section C.2, demonstrate the importance of balancing both objectives to achieve both good calibration and performance.
@@ 7 Conclusion
Limitation and Future Work. Our findings suggest several promising directions for research on LLM calibration. First, determining which regime a model falls into ultimately reduces to understanding whether its accuracy exceeds a certain threshold. This, in turn, depends on two key factors: 1. What is the accuracy threshold for a given neural network architecture? 2. Given such an architecture, which training algorithms lead the model to fall into each regime? Answering these questions requires a deeper theoretical analysis of the properties of transformers, which is currently an open and challenging direction. Second, from a theoretical perspective, determining the constant $C$ in Theorem 4.5 is nontrivial, and we leave it as an open problem for future work. Third, our current study focuses on multiple-choice settings; extending the method to free-form generation is an important direction for future investigation. Fourth, exploring the effectiveness of our approach in scenarios where the LLM is initially poorly calibrated remains an important direction. Finally, exploring the calibration properties of quantized models, black-box API models, and whether low-rank adaptation methods such as LoRA suffer from calibration issues also presents valuable avenues for future research.
@@
Acknowledgments
We would like to thank all the anonymous reviewers for their comments and suggestions. This work was conducted while Ruochen Jin was a visiting scholar at the University of Pennsylvania. This work was supported in part by NIH grants U01CA274576 and P30AG073105, ARPA-H Award D24AC00253, NSF grant DMS-2310679, a Meta Faculty Research Award, Wharton AI for Business, and a PSOM AI2D Seeding Project.
@@ A.1 Preference Optimization
The original DPO method (Rafailov et al., 2023) is to directly optimize the policy without explicitly training the reward function in a supervised manner:
@@ A.2 Other Related Work
Several notable variants of DPO have been developed (Liu et al., 2023; Azar et al., 2024; Chang et al., 2024; Gorbatovski et al., 2024; Rafailov et al., 2024a; Yang et al., 2024). However, recent studies (Li et al., 2023a; Xu et al., 2024a; Tajwar et al., 2024) suggest that DPO is less effective than reward-based RLHF methods for aligning LLMs. Both Li et al. (2023a) and Xu et al. (2024a) attributed this shortcoming to representation misspecification in DPO, which limits its ability to achieve robust alignment compared to reinforcement learning approaches such as PPO. Additionally, the on-policy nature of reward-based fine-tuning helps mitigate distribution shifts between the training data and online responses, thereby enhancing LLM performance (Tajwar et al., 2024).
Nash learning from human feedback (NLHF) has also been proposed as a framework to align LLMs with general preference models (Munos et al., 2023; Liu et al., 2025; Shi et al., 2025). Moreover, Wang et al. (2025) analyzed the convergence behavior of NLHF, showing that it approaches the Nash equilibrium in the last iterate.
Prior work (Guo et al., 2017; Müller et al., 2019) has shown that temperature scaling is generally more effective than label smoothing and other techniques for improving calibration. Zhao et al. (2021) proposed a contextual calibration procedure to improve few-shot performance of language models. Predictive accuracy and calibration trade-off have been studied in general machine learning classification problems (Kumar et al., 2018; Krishnan & Tickoo, 2020; Karandikar et al., 2021; Popordanoska et al., 2022). Additionally, several studies have provided theoretical analyses of the bias in ECE estimators based on uniform-mass binning (UMB) (Gupta et al., 2020; Gupta & Ramdas, 2021). Other related works include Gruber & Buettner (2022) and Sun et al. (2023), which explore the framework of proper calibration errors and minimum-risk recalibration of classifiers.
While temperature scaling (TS) is a strong baseline, our method’s superiority stems from addressing fundamental limitations of post-hoc calibration in LLMs:
  • Direct optimization of calibration: Our approach explicitly minimizes the discrepancy between accuracy and confidence—the definition of ECE—rather than relying on a single scaling parameter. This direct optimization improves calibration while maintaining or enhancing performance. This aligns with previous work on training classifiers with calibration objectives, such as the AvUC loss in Krishnan & Tickoo (2020) and the PAC-Bayes-based objective in Fujisawa & Futami (2024), which demonstrated the superiority of optimization-based strategies over post-hoc methods.
  • Improved generalization under distribution shift: Our approach generalizes better to unseen data and distribution shifts compared to TS, which risks overfitting to specific validation sets. Table 2 clearly demonstrates this advantage in out-domain scenarios—for example, with Olmo2-7B, our CFT method reduces out-domain cw-ECE to 0.0637 (vs. TS’s 0.1196) while improving accuracy to 0.7085 (vs. DPO’s 0.6635). Distribution shifts are common in language tasks, and TS is insufficient to handle such variations.
@@ A.4 Rank Preserving Mapping
We denote the model confidence vector as $\text{Conf}_{\pi}(x)=[c_{1},c_{2},c_{3},c_{4}]$, where $c_{1}\geq c_{2}\geq c_{3}\geq c_{4}$ without loss of generality. Let the calibrated confidence be $\boldsymbol{p}(x)=[c^{\prime}_{1},c^{\prime}_{2},c^{\prime}_{3},c^{\prime}_{4}]$. We set the top-1 calibrated confidence to match the bin accuracy:
$c^{\prime}_{1}=q_{m}.$
To preserve the rank ordering of all four options, we aim to compress the remaining three confidence values $c_{2},c_{3},c_{4}$ into the interval $(0,q_{m})$ while ensuring their sum equals $1-q_{m}$. We apply a nonlinear transformation followed by a linear scaling:
$c^{\prime}_{i}=\alpha\tanh(\gamma c_{i})+\beta,\quad\text{for }i=2,3,4.$
Determining $\gamma$
To control the saturation level of the non-linear function $\tanh(\cdot)$, we select $\gamma$ such that the largest among $c_{2},c_{3},c_{4}$ maps close to 1:
$\tanh\left(\gamma\max\{c_{2},c_{3},c_{4}\}\right)=0.99.$
To ensure numerical stability, we avoid $\tanh^{-1}(1)$ and use the approximation $\text{artanh}(0.99)\approx\ln(3)$. This gives:
$\gamma\max\{c_{2},c_{3},c_{4}\}=\ln(3),\quad\Rightarrow\quad\gamma=\frac{\ln(3)}{\max\{c_{2},c_{3},c_{4}\}}\cdot\frac{1}{1-q_{m}}.$
Solving for $\alpha$ and $\beta$
To satisfy the constraint that $\sum_{i=2}^{4}c^{\prime}_{i}=1-q_{m}$, define:
$T:=\sum_{i=2}^{4}\tanh(\gamma c_{i}).$
Then:
$\sum_{i=2}^{4}c^{\prime}_{i}$ $=\sum_{i=2}^{4}\left(\alpha\tanh(\gamma c_{i})+\beta\right)=\alpha T+3\beta=1-q_{m}.$ (8)
We can now solve for $\alpha$ and $\beta$ in two ways:
(a) Simplified assumption: $\beta=\alpha$
Substitute into Equation (8):
$\alpha T+3\alpha$ $=1-q_{m},$ (9)
$\Rightarrow\quad\alpha(T+3)$ $=1-q_{m},$ (10)
$\Rightarrow\quad\alpha$ $=\frac{1-q_{m}}{T+3},\quad\beta=\alpha.$ (11)
(b) General solution (no assumption):
From Equation (8), isolate $\beta$:
$\alpha T+3\beta$ $=1-q_{m},$ (12)
$\Rightarrow\quad\beta$ $=\frac{1-q_{m}-\alpha T}{3}.$ (13)
Thus, given any $\alpha$, $\beta$ can be directly computed.
Ensuring Rank Preservation
To preserve the ordering $c^{\prime}_{i}<q_{m}$ for $i=2,3,4$, we enforce:
$c^{\prime}_{i}=\alpha\tanh(\gamma c_{i})+\beta<q_{m}.$
Since $\tanh(\gamma c_{i})<1$, a sufficient condition is:
@@
Substitute the simplified solution $\beta=\alpha$:
$\alpha+\beta=2\alpha<q_{m}.$
Then:
$2\alpha<q_{m}\quad$ $\Rightarrow\quad\alpha<\frac{q_{m}}{2},$ (14)
and since $\text{and since }\alpha=\frac{1-q_{m}}{T+3},\quad$ $\Rightarrow\quad\frac{1-q_{m}}{T+3}<\frac{q_{m}}{2}.$ (15)
Multiply both sides of Equation (15) by $(T+3)$:
$2(1-q_{m})$ $<q_{m}(T+3),$ (16)
22 22 $<q_{m}(T+5),$ (17)
$\Rightarrow\quad q_{m}$ $>\frac{2}{T+5}.$ (18)
This inequality gives a general bound on $q_{m}$ for the mapping to preserve the ordering. Noting that $T=\sum_{i=2}^{4}\tanh(\gamma c_{i})<3$, we conservatively estimate $T\leq 3$:
$q_{m}>\frac{2}{T+5}\geq\frac{2}{8}=0.25.$
Conclusion
When $q_{m}>0.25$, the parameters $\alpha$ and $\beta$ can be chosen such that the transformed confidences $c^{\prime}_{i}$ satisfy $c^{\prime}_{i}<q_{m}$ and preserve the original rank ordering. When $q_{m}\leq 0.25$, the non-linear mapping using $\tanh(\gamma\cdot)$ compresses all $c^{\prime}_{i}$ (for $i=2,3,4$) near 1, effectively reflecting high uncertainty with low distinguishability across non-top predictions.
A.5 Convergence of EM Algorithms
The convergence of EM algorithms is well studied; see, for example, Wu (1983). To ensure the convergence of the EM algorithm in our setting, three additional assumptions are required.
First, for the ECE loss,
we assume that $D$ is the cross-entropy loss. In this case,
$\mathcal{L}_{\text{ECE}}=-\mathbb{E}_{\boldsymbol{p}(x)}\left[\log\text{conf}_{\pi}(x)\right].$
Under this assumption, the loss function becomes a negative log-likelihood, which aligns with standard convergence theory for EM algorithms. However, in practice, alternative distance measures may also be used.
Second, the binary choice setting is also required. In this setting, the probability of the false answer is uniquely determined once the probability of the true answer is given. However, in the multiple-choice setting, the generative model assigns probability only to the correct answer, without imposing any constraints on the probabilities of the remaining choices. As a result, these probabilities do not influence the loss function, and the EM algorithm is not uniquely defined in this case. Therefore, the convergence guarantee only applies in the binary setting. In practice, we can use the mapping discussed above to assign probabilities to the remaining choices.
Finally, the optimization is performed over the choice probabilities rather than the neural network parameters, as the convergence of training deep neural networks remains an open problem.
@@ Appendix B Proof of Technical Results
Then, We define $\pi(x)$ as follows. When $x\in\mathcal{S}_{\textnormal{T}}$ or $\mathcal{S}_{\textnormal{F,2}}$,
@@
Then, We define $\pi(x)$ as follows. When $x\in\mathcal{S}_{\textnormal{F}}$ or $\mathcal{S}_{\textnormal{T,2}}$,
@@ C.1 Comprehensive Descriptions for Models, Baselines, Datasets, and Metrics
Models
In our study, we employ four widely-used open-source large language models to investigate the calibration issue and validate the effectiveness of our proposed method. They include
@@
  • Vicuna-7B-v1.5555https://huggingface.co/lmsys/vicuna-7b-v1.5: Vicuna-7B-v1.5 (Chiang et al., 2023) is a chat assistant developed by LMSYS, fine-tuned from Llama 2 using approximately 125,000 user-shared conversations collected from ShareGPT.com. This auto-regressive language model employs the transformer architecture and is designed for research purposes in natural language processing, machine learning, and artificial intelligence. The model has been evaluated using standard benchmarks, human preferences, and LLM-as-a-judge methodologies.
@@
Baselines
We validate the effectiveness of the proposed method by comparing performance among four kinds of models.
@@
Dataset
To evaluate the efficacy of our proposed calibration method, we employ five datasets to conduct comprehensive experiments:
@@
  • AlpacaEval [^12] https://github.com/tatsu-lab/alpaca_eval: Evaluation of instruction-following models (e.g., ChatGPT) typically requires human interaction, which is time-consuming, expensive, and difficult to replicate. AlpacaEval is an LLM-based automatic evaluation framework that is fast, cost-effective, reproducible, and validated against 20K human annotations. It is particularly useful for model development. Although it improves upon prior automatic evaluation pipelines, AlpacaEval still exhibits fundamental limitations, such as a preference for longer outputs. The framework provides the following components:
    • Leaderboard: A leaderboard reporting the performance of common models on the AlpacaEval evaluation set. Caution: Automatic evaluators (e.g., GPT-4) may be biased toward models that produce longer outputs or were fine-tuned on the same base model as the evaluator (e.g., GPT-4).
    • Automatic Evaluator: An evaluator with high agreement with human judgments (validated on 20K annotations). It measures performance by computing the fraction of times a strong LLM (e.g., GPT-4) prefers the outputs from the evaluated model over those from a reference model. The evaluator supports output caching and randomization by default.
    • Toolkit for Building Evaluators: A simple interface for constructing advanced automatic evaluators with features such as caching, batching, multi-annotator support, and statistical analysis (e.g., quality, price, speed, statistical power, bias, and variance).
    • Human Evaluation Data: A dataset of 20K human preferences comparing outputs from a given model and a reference model on the AlpacaFarm evaluation set. This includes 2.5K cross-annotations, where four human annotators rated the same 650 examples.
    • AlpacaEval Dataset: A simplified version of the AlpacaFarm evaluation set, where “instructions” and “inputs” are merged into a single field and reference outputs are extended in length. See the documentation for further details.
  • Arena-Hard [^13] https://github.com/lmarena/arena-hard-auto: Arena-Hard-Auto is an automatic evaluation tool for instruction-tuned LLMs. Among popular open-ended LLM benchmarks, Arena-Hard-Auto demonstrates the highest correlation and separability with LMArena (Chatbot Arena); see the associated paper for details. If you are interested in estimating how well your model might perform on LMArena prior to deployment, we recommend using the latest evaluation set, Arena-Hard-v2.0-Preview.
@@
Metric
Here we introduce three metrics used in our paper: The discrete conf-ECE and cw-ECE, and Win Rate. The discrete conf-ECE and cw-ECE are defined as follows.
Confidence Expected Calibration Error (conf-ECE) is defined as
@@
where $B_{m}$ is the $j$-th bin in terms of max confidence; $|B_{m}|$ denotes the cardinality of the bin; $\mathbb{P}(Y=I|\mathbf{x}\in B_{m})$ and $\mathbb{E}[\pi_{\theta}(\mathbf{y}=I|\mathbf{x})|\mathbf{x}\in B_{m}]$ denote the average prediction of winning class probability and the actual probability that the winning class is the ground truth class.
Class-Wise Expected Calibration Error (cw-ECE) is defined as
@@
We then introduce the detail of the Win Rate metric. Win Rate metric evaluates the model’s “language ability" in two aspects.
The first one is the ability to distinguish between preferred and non-preferred responses within a given dataset. To this end, we utilize a widely-used dataset UltraFeedback  [^15] https://huggingface.co/datasets/argilla/ultrafeedback-binarized-preferences which provides pairs of responses: a chosen_response (preferred) and a reject_response (non-preferred) for each instruction. We then calculate the sequence probabilities of these provided responses under the target model’s distribution. The sequence probability of a response $R=(w_{1},w_{2},\ldots,w_{T})$, where $w_{i}$ denotes the $i$-th token, is computed as:
@@
Here, $P(w_{i}\mid w_{1},w_{2},\ldots,w_{i-1})$ represents the probability of token $w_{i}$ given its preceding context, as determined by the model. For each instruction in the dataset, we calculate the sequence probabilities $P_{\text{chosen}}$ and $P_{\text{reject}}$ for the chosen_response and reject_response, respectively. A “win” is recorded if $P_{\text{chosen}}>P_{\text{reject}}$; otherwise, it is a “loss.”
The second aspect is the quality of the response. We use the questions from two popular datasets AlpacaEval [^16] https://github.com/tatsu-lab/alpaca_eval and Arena-Hard [^17] https://github.com/lmarena/arena-hard-auto and let our model and a baseline model (usually GPT-4/GPT-4o or the DPO/RLHF model that is well aligned with human preference) make two separate responses. We then let a third party large language model (such as GPT-4 or GPT-4o) judge which one is better. A “win” of our model is recorded if the judge thinks its response is better than the baseline; otherwise, it is a “loss.”
The Win Rate is then defined as the proportion of wins over the total number of comparisons:
Win Rate=Number of WinsTotal Number of $\text{Win Rate}=\frac{\text{Number of Wins}}{\text{Total Number of Comparisons}}.$